# Distribution-Guided and Constrained Quantum Machine Unlearning

### v1

@@ 1 Introduction
Quantum Machine Learning (QML) refers to algorithms designed to run on Quantum Computers [7]. With the increased traction of Quantum Machine Learning (QML) in recent years, data privacy, security, and protection remain a persistent issue. [9] design a Membership Inference Attack (MIA) tailored for QML models demonstrating clear evidence of membership leakage from QML models. Their experiments validate Quantum Machine Unlearning (QMU) as a viable mechanism to resolve this issue. Considering the recent traction towards QML, there has been even less work in QMU.
@@
In this work, we introduce a distribution-guided class unlearning framework for variational quantum classifiers that removes the predictive influence of a specified class without retraining from scratch. Our approach formulates unlearning as a constrained optimization problem that suppresses forgotten-class predictions while preserving the model’s behaviour on retained data through soft anchor constraints. Unlike existing unlearning methods that rely on uniform probability suppression or full data deletion, we construct a data-driven forget target distribution that redistributes probability mass according to the semantic similarity inferred from the original model’s outputs. The resulting objective admits a principled Lagrangian interpretation and is optimized directly at the level of quantum circuit parameters using parameter-shift gradients. We demonstrate that this approach achieves selective forgetting through localized parameter updates, maintains retained-class structure, and is effective across datasets of differing complexity, establishing a practical and theoretically grounded unlearning mechanism for near-term quantum machine learning models.
@@ 2 Theoretical Foundations
MU refers to the removal of data and its corresponding influence from a trained machine learning model, without requiring full retraining from scratch. Let $Z$ denote the space of all possible examples, and let $Z^{*}$ denote the set of all finite datasets drawn from $Z$. Given a dataset $D\in Z^{*}$, a learning algorithm $A:Z^{*}\rightarrow H$ maps $D$ to a hypothesis $A(D)$ in the hypothesis space $H$, which represents the set of admissible model parameters.
@@
[4] propose two formal definitions of unlearning, which we summarise below.
@@ 3 Quantum Machine Unlearning
where $D(\cdot\|\cdot)$ denotes a suitable quantum divergence measure. In the context of unlearning, this inequality ensures that the application of a valid quantum operation induces a contractive evolution of model states, preventing increases in distinguishability from the ideal retrained state. Since quantum information cannot be deleted outright, QMU redistributes data-dependent information to the environment while preserving physical consistency.
@@ 4 Methodology
We consider supervised three-class classification with four continuous input features and present a distribution-guided unlearning framework applicable to this setting. The Iris and Covertype datasets are used as representative benchmarks to illustrate the method across tasks of differing complexity. For Covertype, classes 3, 5, and 7 are selected (remapped to 0,1,2), and the original 54-dimensional inputs are reduced to four dimensions via principal component analysis. Each input $x\in\mathbb{R}^{4}$ is min–max scaled to $0,\pi]$ and embedded into a six-qubit variational quantum circuit. Features are encoded on qubits 0–3 using single-qubit rotations $R_{y}(x_{i})$, with pairwise feature interactions introduced via entangling blocks of the form $\mathrm{CX}(a,b)\,R_{z}(x_{a}x_{b})\,\mathrm{CX}(a,b)$ applied to qubit pairs $(0,1)$, $(1,2)$, and $(2,3)$. The feature map is reuploaded twice in a layered feature map–ansatz–feature map–ansatz structure. The ansatz consists of ring entanglement with three repetitions. In each repetition, all six qubits undergo parametrised $R_{y}(\theta)$ and $R_{z}(\theta)$ rotations, followed by a ring of CNOT gates $\mathrm{CX}(q,(q+1)\bmod 6)$. This architecture results in 72 trainable parameters. Classification is performed by measuring Pauli-$Z$ expectation values on qubits 3, 4, and 5, which are treated as logits and passed through a softmax to produce class probabilities. Model parameters are trained by minimizing the mini-batch cross-entropy loss using stochastic gradient-based optimization. Gradients with respect to each circuit parameter are estimated via the parameter-shift rule
@@
The formulation that follows depends only on the model’s predictive distributions and parametrization, and does not rely on dataset-specific structure beyond the assumed number of classes and input features.
@@ 4.1 Distribution-Guided Class Unlearning
Let a trained classifier with parameters $w_{\mathrm{orig}}$ output a $K$-class predictive distribution $p_{w_{\mathrm{orig}}}(y\mid x)$. Given a forgotten class $f$ and a forget set $F$ consisting of training samples with label $f$, our goal is to compute updated parameters $w$ such that: (i) for inputs $x\in F$, the unlearned model assigns negligible probability to the forgotten class, i.e., $p_{w}(f\mid x)$ is suppressed; and (ii) on the retained data, the model’s predictions remain close to the original model (and ideally close to a gold retrained model trained from scratch without $F$). In the following, we operationalize these objectives using a divergence-based forget loss on $F$ and a constraint/regularizer that preserves behaviour on a retained anchor set.
@@ 4.1.4 Optimization and Interpretation
$J(w)=\mathcal{L}_{F}(w)+\alpha\,\mathcal{L}_{A}(w)-\lambda\|w-w_{\mathrm{orig}}\|_{2}^{2},\qquad\alpha,\lambda\geq 0,$ (8)
@@
such that $\frac{1}{|A|}\sum_{x\in A}\mathrm{KL}\!\left(p_{\mathrm{ref}}(\cdot\mid x)\,\|\,p_{w}(\cdot\mid x)\right)\leq\varepsilon,$
@@ 4.3 Results
Unless stated otherwise, all unlearning experiments use $\alpha=1.0$ and $\lambda=0.01$ reflecting equal weighting of the forgetting and preservation terms, together with a mild quadratic anchoring to the original parameters. Targeted unlearning produces a sharp and selective degradation in performance for the forgotten class while largely preserving retained-class behaviour, as shown by the confusion matrices in Figure 4. On Covertype, recall for the forgotten class decreases from $0.633$ to $0.067$, while recall for the two retained classes remains unchanged at $0.467$ and $0.800$, respectively. The mean predicted probability assigned to the forgotten class on its own data drops from $0.4055$ to $0.2715$, indicating a substantial reduction in residual confidence. For Iris, forgotten-class recall decreases from $1.000$ to $0.000$, while recall for class 0 remains at $1.000$ and class 1 decreases modestly from $1.000$ to $0.800$. The corresponding forgotten-class mean probability decreases from $0.4155$ to $0.2098$. Across both datasets, post-unlearning confusion matrices show that errors introduced by unlearning primarily redistribute forgotten samples to a single retained class, consistent with similarity-guided probability reassignment rather than indiscriminate suppression.
Beyond accuracy-based metrics, we evaluate how closely the unlearned model approximates true retraining by comparing its predictive distributions to those of a gold retrained model trained on retained data only. The contraction property of quantum channels in  4 implies that a physically valid unlearning operation cannot increase divergence from the retrained model. We therefore quantify approximate unlearning by directly measuring the remaining divergence to gold retraining.
We compute the Kullback–Leibler (KL) divergence between the gold retrained and unlearned predictive distributions on retained-class test samples. Probabilities are renormalized over retained labels to exclude trivial effects from probability mass assigned to the forgotten class. This ensures that divergence reflects agreement on retained-class behaviour rather than the degree of forgetting itself. This comparison serves as an empirical proxy for distributional equivalence in Definition 2, assessing whether unlearning reproduces the statistical behaviour of retraining on retained data.
On the Covertype dataset, the unlearned model achieves a mean KL divergence of $0.047\pm 0.047$ relative to the gold retrained model, with a median of $0.034$. This low divergence indicates that, despite aggressive suppression of the forgotten class, the unlearned model closely matches the behaviour of a model trained from scratch on retained data. At the same time, the unlearned model assigns a larger fraction of probability mass to the forgotten class than the gold-retrained model on retained samples ($0.295$ vs. $0.195$), reflecting a mild over-forgetting bias that does not substantially affect retained-class predictions. While our experiments focus on class unlearning, extending similarity-guided and constraint-based formulations to instance-level unlearning remains an important direction for future work.
@@
Figure 4: Confusion matrices before (left) and after (right) targeted unlearning. (Left) Covertype dataset with class 2 forgotten. (Right) Iris dataset with class 2 forgotten. In both cases, predictions assigned to the forgotten class are strongly suppressed after unlearning, while errors primarily redistribute to a single retained class.
@@
5 Conclusion
We introduced a distribution-guided and constrained framework for quantum machine unlearning that selectively suppresses forgotten-class influence while preserving retained behaviour. Experiments on variational quantum classifiers demonstrate sharp forgetting, minimal degradation on retained classes, and close alignment with gold retraining baselines. From a quantum information perspective, our approach shows how the abstract idea of contractive quantum forgetting can be realized in practice, by using distribution-guided targets and anchor constraints to achieve controlled, measurable unlearning in variational quantum classifiers as demonstrated on Iris and Covertype as representative examples.
@@ Appendix B Proof of Theorem 1
for a constant $C$ independent of $w$.
Substituting this expression into (8) and discarding constants independent of $w$ gives

### v2

@@ 1 Introduction
Quantum Machine Learning (QML) refers to algorithms designed to run on Quantum Computers [7]. With the increased traction of Quantum Machine Learning (QML) in recent years, data privacy, security, and protection remain a persistent issue. [9] design a Membership Inference Attack (MIA) tailored for QML models, demonstrating clear evidence of membership leakage from QML models. Their experiments validate Quantum Machine Unlearning (QMU) as a viable mechanism to resolve this issue. Considering the recent traction towards QML, there has been even less work in QMU.
@@
In this work, we introduce a distribution-guided class unlearning framework for variational quantum classifiers that remove the predictive influence of a specified class without retraining from scratch. Our approach formulates unlearning as a constrained optimization problem that suppresses forgotten-class predictions while preserving the model’s behaviour on retained data through soft anchor constraints. Unlike existing unlearning methods that rely on uniform probability suppression or full data deletion, we construct a data-driven forget target distribution that redistributes probability mass according to the semantic similarity inferred from the original model’s outputs. The resulting objective admits a principled Lagrangian interpretation and is optimized directly at the level of quantum circuit parameters using parameter-shift gradients. We demonstrate that this approach achieves selective forgetting through localized parameter updates, maintains retained-class structure, and is effective across datasets of differing complexity, establishing a practical and theoretically grounded unlearning mechanism for near-term quantum machine learning models.
@@ 2 Theoretical Foundations
Machine unlearning refers to the removal of data and its corresponding influence from a trained machine learning model, without requiring full retraining from scratch. Let $Z$ denote the space of all possible examples, and let $Z^{*}$ denote the set of all finite datasets drawn from $Z$. Given a dataset $D\in Z^{*}$, a learning algorithm $A:Z^{*}\rightarrow H$ maps $D$ to a hypothesis $A(D)$ in the hypothesis space $H$, which represents the set of admissible model parameters.
@@
[4] proposes two formal definitions of unlearning, which we summarise below.
@@ 3 Quantum Machine Unlearning
where $D(\cdot\|\cdot)$ denotes a suitable quantum divergence measure. In the context of unlearning, this inequality motivates viewing forgetting as a contractive reduction of data-dependent distinguishability, consistent with physically admissible quantum evolutions. Since quantum information cannot be deleted outright, QMU redistributes data-dependent information to the environment while preserving physical consistency. Accordingly, for variational quantum classifiers, contractivity is reflected in reduced distinguishability at the level of observable predictive distributions.
@@ 4 Methodology
We consider supervised three-class classification with four continuous input features and present a distribution-guided unlearning framework applicable to this setting. We instantiate the framework on two representative classification problems — one low-complexity (Iris) and one higher-complexity (Covertype) — to illustrate how the same unlearning objective behaves across regimes. For Covertype, classes 3, 5, and 7 are selected (remapped to 0,1,2), and the original 54-dimensional inputs are reduced to four dimensions via principal component analysis. Each input, in this case $x\in\mathbb{R}^{4}$, is min–max scaled to $0,\pi]$ and embedded into a six-qubit variational quantum circuit. Features are encoded on qubits 0–3 using single-qubit rotations $R_{y}(x_{i})$, with pairwise feature interactions introduced via entangling blocks of the form $\mathrm{CX}(a,b)\,R_{z}(x_{a}x_{b})\,\mathrm{CX}(a,b)$ applied to qubit pairs $(0,1)$, $(1,2)$, and $(2,3)$. The feature map is reuploaded twice in a layered feature map–ansatz–feature map–ansatz structure. The ansatz consists of ring entanglement with three repetitions. In each repetition, all six qubits undergo parametrised $R_{y}(\theta)$ and $R_{z}(\theta)$ rotations, followed by a ring of CNOT gates $\mathrm{CX}(q,(q+1)\bmod 6)$. This architecture results in 72 trainable parameters. Classification is performed by measuring Pauli-$Z$ expectation values on qubits 3, 4, and 5, which are treated as logits and passed through a softmax to produce class probabilities. Model parameters are trained by minimizing the mini-batch cross-entropy loss using stochastic gradient-based optimization. Gradients with respect to each circuit parameter are estimated via the parameter-shift rule
@@
The formulation that follows depends only on the model’s predictive distributions and parameter space, and does not rely on dataset-specific structure beyond a fixed classification setting. The proposed unlearning objective is defined at the level of output distributions and model parameters, and is independent of the particular dataset, feature embedding, or circuit architecture used to instantiate the classifier.
@@ 4.1 Distribution-Guided Class Unlearning
Let a trained classifier with parameters $w_{\mathrm{orig}}$ output a $K$-class predictive distribution $p_{w_{\mathrm{orig}}}(y\mid x)$. Given a forgotten class $f$ and a forget set $F$ consisting of training samples with label $f$, our goal is to compute updated parameters $w$ such that: (i) for inputs $x\in F$, the unlearned model assigns negligible probability to the forgotten class, i.e., $p_{w}(f\mid x)$ is suppressed; and (ii) on the retained data, the model’s predictions remain close to the original model (and ideally close to a gold retrained model trained from scratch without $F$). In the following, we operationalize these objectives using a divergence-based forget loss on $F$ and a constraint that preserves behaviour on a retained anchor set.
@@ 4.1.4 Optimization and Interpretation
$J(w)=\mathcal{L}_{F}(w)+\alpha\,\mathcal{L}_{A}(w)-\lambda\|w-w_{\mathrm{orig}}\|_{2}^{2},\quad\alpha,\lambda\geq 0,$ (8)
@@
[…]
@@ 4.3 Results
Unless stated otherwise, all unlearning experiments use $\alpha=1.0$ and $\lambda=0.01$ reflecting equal weighting of the forgetting and preservation terms, together with a mild quadratic anchoring to the original parameters. Targeted unlearning produces a sharp and selective degradation in performance for the forgotten class, while largely preserving retained-class behaviour, as shown in Figures 4 and 5. On Covertype, recall for the forgotten class decreases from $0.633$ to $0.067$, while recall for the two retained classes remains unchanged at $0.467$ and $0.800$, respectively. The mean predicted probability assigned to the forgotten class on its own data drops from $0.4055$ to $0.2715$, indicating a substantial reduction in residual confidence. For Iris, forgotten-class recall decreases from $1.000$ to $0.000$, while recall for class 0 remains at $1.000$ and class 1 decreases modestly from $1.000$ to $0.800$. The corresponding forgotten-class mean probability decreases from $0.4155$ to $0.2098$. Across both datasets, post-unlearning confusion matrices show that errors introduced by unlearning primarily redistribute forgotten samples to a single retained class, consistent with similarity-guided probability reassignment rather than indiscriminate suppression.
@@
Figure 4: Confusion matrices before and after targeted unlearning on the Covertype dataset with class 2 forgotten. Predictions assigned to the forgotten class are strongly suppressed after unlearning, with errors primarily redistributing to a single retained class.
Figure 5: Confusion matrices before and after targeted unlearning on the Iris dataset with class 2 forgotten. Forgotten-class predictions are eliminated, while retained classes largely preserve pre-unlearning behaviour.
Beyond accuracy-based metrics, we evaluate how closely the unlearned model approximates true retraining by comparing its predictive distributions to those of a gold retrained model trained on retained data only. We compute the Kullback–Leibler (KL) divergence between the gold retrained and unlearned predictive distributions on retained-class test samples. Probabilities are renormalized over retained labels to exclude trivial effects from probability mass assigned to the forgotten class. This ensures that divergence reflects agreement on retained-class behaviour rather than the degree of forgetting itself. This comparison serves as an empirical proxy for distributional equivalence in Definition 2, assessing whether unlearning reproduces the statistical behaviour of retraining on retained data.
On the Covertype dataset, the unlearned model achieves a mean KL divergence of $0.047\pm 0.047$ relative to the gold retrained model, with a median of $0.034$. This low divergence indicates that, despite aggressive suppression of the forgotten class, the unlearned model closely matches the behaviour of a model trained from scratch on retained data. At the same time, the unlearned model assigns a larger fraction of probability mass to the forgotten class than the gold-retrained model on retained samples ($0.295$ vs. $0.195$), reflecting a mild under-learning that does not substantially affect retained-class predictions. While our experiments focus on class unlearning, extending similarity-guided and constraint-based formulations to instance-level unlearning remains an important direction for future work.
@@
5 Conclusions
We introduced a generalised four-feature three-class distribution-guided and constrained framework for quantum machine unlearning that selectively suppresses forgotten-class influence while preserving retained behaviour. Experiments on variational quantum classifiers demonstrate sharp forgetting, minimal degradation on retained classes, and close alignment with gold retraining baselines. From a quantum information perspective, our framework is motivated by physical admissibility constraints on unlearning operations. In practice, we realize these principles through constrained variational optimization, with unlearning quality evaluated empirically via alignment to gold retraining.
@@ Appendix B Proof of Theorem 1
for a constant $C$ independent of $w$. Substituting this expression into (8) and discarding constants independent of $w$ gives
@@
Appendix C Ablation Studies
This appendix investigates the sensitivity of the proposed distribution-guided quantum machine unlearning framework to its principal design choices and hyperparameters. Given that the Iris dataset is a simple, low-complexity classification task, all ablations are conducted on the Covertype dataset using the variational quantum classifier described in Section 4.1, unless stated otherwise. Unless varied explicitly, the default unlearning parameters are $\alpha=1.0$, $\lambda=0.01$, and $\beta=1.0$. To avoid repeating the class unlearning discussed in the previous discussions, we exclude unlearning in class 2 in these experiments.
Throughout this appendix, forgetting effectiveness is quantified by the mean predicted probability assigned to the forgotten class on its own samples, while retained utility is measured by test accuracy on all classes.
C.1 Class-wise forgetting robustness
We first evaluate whether the proposed method depends on the specific choice of forgotten class. Using the default unlearning configuration, we apply class-level unlearning to each non-trivial class in turn and report the resulting forgetting strength and retained accuracy.
Table 3: Class-wise unlearning performance on Covertype. Results are reported for non-trivial forgotten classes. Lower forgotten-class probability indicates stronger forgetting.
Forgotten Class Test Acc. (Before) Test Acc. (After) $p_{f}$ (After)
0 0.633 0.411 0.260
1 0.633 0.411 0.261
Across forgotten classes, the method consistently suppresses confidence assigned to the forgotten class while maintaining comparable retained accuracy. This indicates that the unlearning behaviour is not specific to a single class choice and is robust across non-trivial forgetting scenarios.
C.2 Sensitivity to similarity sharpness $\beta$
The distribution-guided target distribution in Eq. (6) depends on the sharpness parameter $\beta$, which controls how strongly the probability mass is concentrated on the most semantically similar retained class. We sweep $\beta\in\{0.25,0.5,0.75,1.0\}$ while keeping all other parameters fixed.
Table 4: Effect of varying the similarity sharpness parameter $\beta$ on forgetting and retention.
$\beta$ Test Acc. (After) $p_{f}$ (After)
0.25 0.400 0.269
0.50 0.411 0.267
0.75 0.422 0.264
1.00 0.411 0.261
Performance varies smoothly with $\beta$, with no sharp degradation in either forgetting or retaining accuracy. This indicates that the similarity-guided target construction is robust to moderate changes in sharpness, and does not require precise tuning.
C.3 Anchor set size sensitivity
We next examine how the size of the retained anchor set affects unlearning performance. The anchor set is subsampled in a stratified manner to preserve class balance, using fractions $\{10\%,25\%,50\%,100\%\}$ of the original anchor data.
Table 5: Effect of anchor set size on unlearning performance.
Anchor Fraction Test Acc. (After) $p_{f}$ (After)
0.10 0.356 0.250
0.25 0.411 0.260
0.50 0.422 0.260
1.00 0.411 0.261
Retention performance improves rapidly as the anchor fraction increases from $10\%$ to $25\%$ and stabilizes thereafter, while forgetting effectiveness remains largely unchanged. This suggests that only a modest retained anchor subset is sufficient to preserve model behaviour, highlighting the practical efficiency of the proposed constraint.
C.4 Effect of anchor-preservation weight $\alpha$
The coefficient $\alpha$ in Eq. (7) controls the strength of the anchor-preservation term. To assess its role, we vary $\alpha\in\{0,1,2\}$ while keeping all other parameters fixed.
Table 6: Effect of varying $\alpha$ on forgetting and retention.
$\alpha$ Test Acc. (After) $p_{f}$ (After)
0 0.300 0.214
1 0.411 0.261
2 0.400 0.270
Removing the anchor constraint ($\alpha=0$) leads to aggressive forgetting but substantial degradation of retained accuracy. Increasing $\alpha$ moderates forgetting while improving retention, consistent with the Lagrangian interpretation of Theorem 1.
C.5 Effect of parameter anchoring $\lambda$
Finally, we study the effect of the quadratic parameter anchoring term in Eq. (7) by varying $\lambda\in\{0,0.01,0.1\}$.
Table 7: Effect of varying the parameter anchoring coefficient $\lambda$.
$\lambda$ Test Acc. (After) $p_{f}$ (After)
0 0.411 0.251
0.01 0.411 0.261
0.10 0.433 0.334
Larger values of $\lambda$ improve retention at the cost of weaker forgetting, while $\lambda=0$ retains effective forgetting but provides less stability. This confirms that parameter anchoring acts as a secondary stabilization mechanism, complementing the anchor distribution constraint rather than replacing it.
Summary
Across all ablations, the proposed distribution-guided and constrained unlearning framework exhibits stable and interpretable behaviour. The method is robust to the choice of forgotten class, insensitive to moderate changes in the similarity sharpness parameter, and requires only a modest anchor subset to preserve retained behaviour. The effects of $\alpha$ and $\lambda$ align closely with the Lagrangian interpretation of the objective, providing empirical validation of the theoretical formulation in Section 4.1.